# Improving robustness to corruptions with multiplicative weight perturbations

### v1

@@ Abstract
Deep neural networks (DNNs) excel on clean images but struggle with corrupted ones. Incorporating specific corruptions into the data augmentation pipeline can improve robustness to those corruptions but may harm performance on clean images and other types of distortion. In this paper, we introduce an alternative approach that improves the robustness of DNNs to a wide range of corruptions without compromising accuracy on clean images. We first demonstrate that input perturbations can be mimicked by multiplicative perturbations in the weight space. Leveraging this, we propose Data Augmentation via Multiplicative Perturbation (DAMP), a training method that optimizes DNNs under random multiplicative weight perturbations. We also examine the recently proposed Adaptive Sharpness-Aware Minimization (ASAM) and show that it optimizes DNNs under adversarial multiplicative weight perturbations. Experiments on image classification datasets (CIFAR-10/100, TinyImageNet and ImageNet) and neural network architectures (ResNet50, ViT-S/16) show that DAMP enhances model generalization performance in the presence of corruptions across different settings. Notably, DAMP is able to train a ViT-S/16 on ImageNet from scratch, reaching the top-1 error of $23.7\%$ which is comparable to ResNet50 without extensive data augmentations.
@@ 1 Introduction
In this work, we show that simply perturbing weights with random multiplicative noises during training can significantly improve robustness to a wide range of corruptions. Our contributions are as follows:
  • We show in Section 2 that the effects of input corruptions can be simulated during training via multiplicative weight perturbations.
  • From this insight, we propose a new training algorithm called Data Augmentation via Multiplicative Perturbations (DAMP) which perturbs weights using random multiplicative Gaussian noise during training while having the same training cost as standard SGD.
@@
  • Notably, we demonstrate that DAMP can train a Vision Transformer (ViT) (Dosovitskiy et al., 2021) from scratch on ImageNet, achieving similar accuracy to a ResNet50 (He et al., 2016a) in 200 epochs with only basic Inception-style preprocessing (Szegedy et al., 2016). This is significant as ViT typically requires advanced training methods or sophisticated data augmentation to match ResNet50’s performance when being trained on ImageNet from scratch (Chen et al., 2022; Beyer et al., 2022).
@@ 2 Data Augmentation via Multiplicative Perturbations
In this section, we show that one can simulate the effect of input corruptions during training via multiplicative weight perturbations. Given a training data set $\mathcal{S}=\{(\mathbf{x}_{k},y_{k})\}_{k=1}^{N}\subseteq\mathcal{X}\times\mathcal{Y}$ drawn i.i.d. from the data distribution $\mathcal{D}$, we seek to learn a model that generalizes well on both clean and corrupted inputs. We denote $\mathcal{G}$ as a set of functions whose each member $\mathbf{g}:\mathcal{X}\rightarrow\mathcal{X}$ represents an input corruption. That is, for each $\mathbf{x}\in\mathcal{X}$, $\mathbf{g}(\mathbf{x})$ is a corrupted version of $\mathbf{x}$ [^1] For instance, if $\mathbf{x}$ is a clean image then $\mathbf{g}(\mathbf{x})$ could be $\mathbf{x}$ corrupted by Gaussian noise. We define $\mathbf{g}(\mathcal{S})\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\{(\mathbf{g}(\mathbf{x}_{k}),y_{k})\}_{k=1}^{N}$ as the training set corrupted by $\mathbf{g}$. We consider a fully-connected neural network $\mathbf{f}(\mathbf{x};\boldsymbol{\omega})$ of depth $H$ parameterized by $\boldsymbol{\omega}=\{\mathbf{W}^{(h)}\}_{h=1}^{H}\in\mathcal{W}$, which we define recursively as follows:
$\mathbf{f}^{(0)}(\mathbf{x})$ =$\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\mathbf{x},$ (1)
$\mathbf{z}^{(h)}(\mathbf{x})$ =$\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\mathbf{W}^{(h)}\mathbf{f}^{(h-1)}(\mathbf{x}),\qquad\mathbf{f}^{(h)}(\mathbf{x})\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\boldsymbol{\sigma}^{(h)}(\mathbf{z}^{(h)}(\mathbf{x})),\qquad\forall h=1,\dots,H$ (2)
$\mathbf{f}(\mathbf{x};\boldsymbol{\omega})$ =$\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\mathbf{f}^{(H)}(\mathbf{x}).$ (3)
where $\boldsymbol{\sigma}^{(h)}$ is the non-linear activation of layer $h$. Given a per-sample loss $\ell:\mathcal{W}\times\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}_{+}$, the training loss is defined as the average loss over the samples $\mathcal{L}(\boldsymbol{\omega};\mathcal{S})\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\frac{1}{N}\sum_{k=1}^{N}\ell(\boldsymbol{\omega},\mathbf{x}_{k},y_{k})$. Our goal is to find $\boldsymbol{\omega}$ which minimizes:
$\mathcal{L}(\boldsymbol{\omega};\mathcal{G}(\mathcal{S}))\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\mathbb{E}_{\mathbf{g}\sim\mathcal{G}}[\mathcal{L}(\boldsymbol{\omega};\mathbf{g}(\mathcal{S}))]$ (4)
Now we show that the effect of each corruption $\mathbf{g}$ can be simulated during training with multiplicative weight perturbations. To this end, we study how input corruption affects the training loss. We use the following two assumptions for our analysis:
@@
The gradient of the per-sample loss with respect to the input $\nabla_{\mathbf{x}}\ell(\boldsymbol{\omega},\mathbf{x},y)$ is Lipschitz continuous.
Assumption 1 means that the corruption $\mathbf{g}$ does not substantially alter the input $\mathbf{x}$, which is true for images since the values of image pixels are bounded between 00 and $1$. Assumption 2 allows us to define a quadratic bound of the loss function using a second-order Taylor expansion. We first notice that the corruption $\mathbf{g}$ causes a shift $\boldsymbol{\delta}_{\mathbf{g}}\mathbf{f}^{(0)}(\mathbf{x})\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\mathbf{g}(\mathbf{x})-\mathbf{x}$ in the input $\mathbf{x}$ which leads to shifts in the output of each layer during the forward pass:
$\underbrace{\boldsymbol{\delta}_{\mathbf{g}}\mathbf{f}^{(h)}(\mathbf{x})}_{\text{shift}}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\underbrace{\mathbf{f}^{(h)}(\mathbf{g}(\mathbf{x}))}_{\text{corrupted output}}-\underbrace{\mathbf{f}^{(h)}(\mathbf{x})}_{\text{clean output}}$ (5)
This will eventually cause a shift in the per-sample loss $\boldsymbol{\delta}_{\mathbf{g}}\ell(\boldsymbol{\omega},\mathbf{x},y)\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\ell(\boldsymbol{\omega},\mathbf{g}(\mathbf{x}),y)-\ell(\boldsymbol{\omega},\mathbf{x},y)$. Using Assumptions 1 and 2, we prove the following lemma which characterizes the connection between $\boldsymbol{\delta}_{\mathbf{g}}\ell(\boldsymbol{\omega},\mathbf{x},y)$ and $\boldsymbol{\delta}_{\mathbf{g}}\mathbf{f}^{(h)}(\mathbf{x})$:
@@
$\boldsymbol{\delta}_{\mathbf{g}}\ell(\boldsymbol{\omega},\mathbf{x},y)\leq{\left\langle\nabla_{\mathbf{z}^{(h+1)}}\ell(\boldsymbol{\omega},\mathbf{x},y)\otimes\boldsymbol{\delta}_{\mathbf{g}}\mathbf{f}^{(h)}(\mathbf{x}),\mathbf{W}^{(h+1)}\right\rangle}_{F}+\frac{C_{\mathbf{g}}^{(h)}(\mathbf{x})}{2}||\mathbf{W}^{(h)}||_{F}^{2}$ (6)
Here $\otimes$ denotes the outer product of two vectors, $\langle\cdot,\cdot\rangle_{F}$ denotes the Frobenius inner product of two matrices of the same dimension, $||\cdot||_{F}$ is the Frobenius norm, and $\nabla_{\mathbf{z}^{(h)}}\ell(\boldsymbol{\omega},\mathbf{x},y)$ is the Jacobian of the per-sample loss with respect to the pre-activation output $\mathbf{z}^{(h)}(\mathbf{x})$ at layer $h$. The proof of Lemma 1 is provided in Appendix A. Using Lemma 1, we prove the following theorem, which bounds the training loss in the presence of corruptions using the training loss under multiplicative perturbations in the weight space:
@@
For a function $\mathbf{g}:\mathcal{X}\rightarrow\mathcal{X}$ satisfying Assumption 1 and a loss function $\mathcal{L}$ satisfying Assumption 2, there exists $\boldsymbol{\xi}_{g}\in\mathcal{W}$ and $C_{\mathbf{g}}>0$ such that:
$\mathcal{L}(\boldsymbol{\omega};\mathbf{g}(\mathcal{S}))\leq\mathcal{L}(\boldsymbol{\omega}\circ\boldsymbol{\xi}_{g};\mathcal{S})+\frac{C_{\mathbf{g}}}{2}||\boldsymbol{\omega}||_{F}^{2}$ (7)
where $\circ$ denotes the Hadamard product.
Proof.
From Lemma 1, we have for all $h=0,\dots,H-1$:
$\leq\mathcal{L}(\boldsymbol{\omega};\mathcal{S})+\frac{1}{N}\sum_{k=1}^{N}{\left\langle\nabla_{\mathbf{z}^{(h+1)}}\ell(\boldsymbol{\omega},\mathbf{x}_{k},y_{k})\otimes\boldsymbol{\delta}_{\mathbf{g}}\mathbf{f}^{(h)}(\mathbf{x}_{k}),\mathbf{W}^{(h+1)}\right\rangle}_{F}+\frac{\hat{C}_{\mathbf{g}}^{(h)}}{2}||\mathbf{W}^{(h)}||_{F}^{2}$ (9)
$\mathcal{L}(\boldsymbol{\omega};\mathbf{g}(\mathcal{S}))\leq\mathcal{L}(\boldsymbol{\omega};\mathcal{S})+\frac{1}{H}\sum_{h=1}^{H}\frac{1}{N}\sum_{k=1}^{N}{\left\langle\nabla_{\mathbf{z}^{(h)}}\ell(\boldsymbol{\omega},\mathbf{x}_{k},y_{k})\otimes\boldsymbol{\delta}_{\mathbf{g}}\mathbf{f}^{(h-1)}(\mathbf{x}_{k}),\mathbf{W}^{(h)}\right\rangle}_{F}\\ +\frac{C_{\mathbf{g}}}{2}||\boldsymbol{\omega}||_{F}^{2}$ (10)
$\mathcal{L}(\boldsymbol{\omega};\mathcal{S})+\frac{1}{H}\sum_{h=1}^{H}{\left\langle\frac{1}{N}\sum_{k=1}^{N}\nabla_{\mathbf{z}^{(h)}}\ell(\boldsymbol{\omega},\mathbf{x}_{k},y_{k})\otimes\boldsymbol{\delta}_{\mathbf{g}}\mathbf{f}^{(h-1)}(\mathbf{x}_{k}),\mathbf{W}^{(h)}\right\rangle}_{F}+\frac{C_{\mathbf{g}}}{2}||\boldsymbol{\omega}||_{F}^{2}$ (11)
$\left[\boldsymbol{\xi}^{(h)}(\mathbf{g})\right]_{i,j}=\frac{1}{H}\frac{\left[\sum_{k=1}^{N}\nabla_{\mathbf{z}^{(h)}}\ell(\boldsymbol{\omega},\mathbf{x}_{k},y_{k})\otimes\boldsymbol{\delta}_{\mathbf{g}}\mathbf{f}^{(h-1)}(\mathbf{x}_{k})\right]_{i,j}}{\left[\sum_{k=1}^{N}\nabla_{\mathbf{z}^{(h)}}\ell(\boldsymbol{\omega},\mathbf{x}_{k},y_{k})\otimes\mathbf{f}^{(h-1)}(\mathbf{x}_{k})\right]_{i,j}}$ (14)
Theorem 1 provides an upper bound to the target loss in Eq. 4:
$\mathcal{L}(\boldsymbol{\omega};\mathcal{G}(\mathcal{S}))\leq\mathbb{E}_{\mathbf{g}\sim\mathcal{G}}\left[\mathcal{L}(\boldsymbol{\omega}\circ(1+\boldsymbol{\xi}(\mathbf{g}));\mathcal{S})+\frac{C_{\mathbf{g}}}{2}||\boldsymbol{\omega}||_{F}^{2}\right]$ (15)
This bound suggests that in a scenario where we do not have any information regarding the set of corruptions $\mathcal{G}$, we can simulate $\mathcal{G}$ using random multiplicative weight perturbations, leading to the following proxy loss:
$\mathcal{L}_{\boldsymbol{\Xi}}(\boldsymbol{\omega};\mathcal{S})\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\mathbb{E}_{\boldsymbol{\xi}\sim\boldsymbol{\Xi}}\left[\mathcal{L}(\boldsymbol{\omega}\circ\boldsymbol{\xi};\mathcal{S})\right]+\frac{\lambda}{2}||\boldsymbol{\omega}||_{F}^{2}$ (16)
We note that the second term in Eq. 16 is the $L_{2}$-regularization commonly used in optimizing DNNs. Based on this proxy loss, we propose Algorithm 1 which minimizes the objective function in Eq. 16 when $\boldsymbol{\Xi}$ is an isotropic Gaussian distribution $\mathcal{N}(\boldsymbol{1},\sigma^{2}\mathbf{I})$. We call this algorithm Data Augmentation via Multiplicative Perturbations (DAMP), as it uses random multiplicative weight perturbations during training to simulate input corruptions, which can be viewed as data augmentations.
@@
5:     Draw a mini-batch =$\mathcal{B}=\{\mathbf{x}_{b},y_{b}\}_{b=1}^{B}\sim\mathcal{S}$.
@@
The standard method to calculate the expected loss in Eq. 16, which lacks a closed-form solution, is the Monte Carlo (MC) approximation. However, the training cost of this approach scales linearly with the number of MC samples. To match the training cost of standard SGD, Algorithm 1 divides each data batch into $M$ equal sub-batches (6) and calculates the loss on each sub-batch with different multiplicative noises from the noise distribution $\boldsymbol{\Xi}$ (79). The final gradient is obtained by averaging the sub-batch gradients (11). Algorithm 1 is thus suitable for data parallelism in multi-GPU training, where the data batch is evenly distributed across $M>1$ GPUs. Compared to SGD, Algorithm 1 requires only two additional operations: generating Gaussian samples and point-wise multiplication, both of which have negligible computational costs. In our experiments, we found that both SGD and DAMP had similar training times.
@@
In this section, we demonstrate that optimizing DNNs with adversarial multiplicative weight perturbations follows a similar update rule to Adaptive Sharpness-Aware Minimization (ASAM) (Kwon et al., 2021). We first provide a brief description of ASAM and its predecessor Sharpness-Aware Minimization (SAM) (Foret et al., 2021):

### v2

@@ Abstract
Deep neural networks (DNNs) excel on clean images but struggle with corrupted ones. Incorporating specific corruptions into the data augmentation pipeline can improve robustness to those corruptions but may harm performance on clean images and other types of distortion. In this paper, we introduce an alternative approach that improves the robustness of DNNs to a wide range of corruptions without compromising accuracy on clean images. We first demonstrate that input perturbations can be mimicked by multiplicative perturbations in the weight space. Leveraging this, we propose Data Augmentation via Multiplicative Perturbation (DAMP), a training method that optimizes DNNs under random multiplicative weight perturbations. We also examine the recently proposed Adaptive Sharpness-Aware Minimization (ASAM) and show that it optimizes DNNs under adversarial multiplicative weight perturbations. Experiments on image classification datasets (CIFAR-10/100, TinyImageNet and ImageNet) and neural network architectures (ResNet50, ViT-S/16, ViT-B/16) show that DAMP enhances model generalization performance in the presence of corruptions across different settings. Notably, DAMP is able to train a ViT-S/16 on ImageNet from scratch, reaching the top-1 error of $23.7\%$ which is comparable to ResNet50 without extensive data augmentations. [^1] Our code is available at https://github.com/trungtrinh44/DAMP
@@ 1 Introduction
(a) $z=\mathbf{w}^{\top}(\mathbf{x}+{\color[rgb]{0.72265625,0.328125,0.3125}\boldsymbol{\epsilon}})$
(b) $z=(\mathbf{w}\circ{\color[rgb]{0.72265625,0.328125,0.3125}(1+\boldsymbol{\epsilon}/\mathbf{x})})^{\top}\mathbf{x}$
(c) $z=(\mathbf{w}\circ{\color[rgb]{0.2890625,0.4765625,0.4453125}\boldsymbol{\xi}})^{\top}\mathbf{x},\,\,{\color[rgb]{0.2890625,0.4765625,0.4453125}\boldsymbol{\xi}}\sim p({\color[rgb]{0.2890625,0.4765625,0.4453125}\boldsymbol{\xi}})$
Figure 1: Depictions of a pre-activation neuron $z=\mathbf{w}^{\top}\mathbf{x}$ in the presence of (a) covariate shift ${\color[rgb]{0.72265625,0.328125,0.3125}\boldsymbol{\epsilon}}$, (b) a multiplicative weight perturbation (MWP) equivalent to ${\color[rgb]{0.72265625,0.328125,0.3125}\boldsymbol{\epsilon}}$, and (c) random MWPs ${\color[rgb]{0.2890625,0.4765625,0.4453125}\boldsymbol{\xi}}$. $\circ$ denotes the Hadamard product. Figs. (a) and (b) show that for a covariate shift ${\color[rgb]{0.72265625,0.328125,0.3125}\boldsymbol{\epsilon}}$, one can always find an equivalent MWP. From this intuition, we propose to inject random MWPs ${\color[rgb]{0.2890625,0.4765625,0.4453125}\boldsymbol{\xi}}$ to the forward pass during training as shown in Fig. (c) to robustify a DNN to covariate shift.
@@
In this work, we show that simply perturbing weights with multiplicative random variables during training can significantly improve robustness to a wide range of corruptions. Our contributions are as follows:
  • We show in Section 2 and Fig. 1 that the effects of input corruptions can be simulated during training via multiplicative weight perturbations.
  • From this insight, we propose a new training algorithm called Data Augmentation via Multiplicative Perturbations (DAMP) which perturbs weights using multiplicative Gaussian random variables during training while having the same training cost as standard SGD.
@@
  • Notably, we demonstrate that DAMP can train a Vision Transformer (ViT) (Dosovitskiy et al., 2021) from scratch on ImageNet, achieving similar accuracy to a ResNet50 (He et al., 2016a) in 200 epochs with only basic Inception-style preprocessing (Szegedy et al., 2016). This is significant as ViT typically requires advanced training methods or sophisticated data augmentation to match ResNet50’s performance when being trained on ImageNet from scratch (Chen et al., 2022; Beyer et al., 2022). We also show that DAMP can be combined with modern augmentation techniques such as MixUp (Zhang et al., 2018) and RandAugment (Cubuk et al., 2020) to further improve robustness of neural networks.
@@ 2 Data Augmentation via Multiplicative Perturbations
In this section, we demonstrate the equivalence between input corruptions and multiplicative weight perturbations (MWPs), as shown in Fig. 1, motivating the use of MWPs for data augmentation.
2.1 Problem setting
Given a training data set $\mathcal{S}=\{(\mathbf{x}_{k},y_{k})\}_{k=1}^{N}\subseteq\mathcal{X}\times\mathcal{Y}$ drawn i.i.d. from the data distribution $\mathcal{D}$, we seek to learn a model that generalizes well on both clean and corrupted inputs. We denote $\mathcal{G}$ as a set of functions whose each member $\mathbf{g}:\mathcal{X}\rightarrow\mathcal{X}$ represents an input corruption. That is, for each $\mathbf{x}\in\mathcal{X}$, $\mathbf{g}(\mathbf{x})$ is a corrupted version of $\mathbf{x}$ [^2] For instance, if $\mathbf{x}$ is a clean image then $\mathbf{g}(\mathbf{x})$ could be $\mathbf{x}$ corrupted by Gaussian noise. We define $\mathbf{g}(\mathcal{S})\vcentcolon=\{(\mathbf{g}(\mathbf{x}_{k}),y_{k})\}_{k=1}^{N}$ as the training set corrupted by $\mathbf{g}$. We consider a DNN $\mathbf{f}:\mathcal{X}\rightarrow\mathcal{Y}$ parameterized by $\boldsymbol{\omega}\in\mathcal{W}$. Given a per-sample loss $\ell:\mathcal{W}\times\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}_{+}$, the training loss is defined as the average loss over the samples $\mathcal{L}(\boldsymbol{\omega};\mathcal{S})\vcentcolon=\frac{1}{N}\sum_{k=1}^{N}\ell(\boldsymbol{\omega},\mathbf{x}_{k},y_{k})$. Our goal is to find $\boldsymbol{\omega}$ which minimizes:
$\mathcal{L}(\boldsymbol{\omega};\mathcal{G}(\mathcal{S}))\vcentcolon=\mathbb{E}_{\mathbf{g}\sim\mathcal{G}}[\mathcal{L}(\boldsymbol{\omega};\mathbf{g}(\mathcal{S}))]$ (1)
without knowing exactly the types of corruption contained in $\mathcal{G}$. This problem is crucial for the reliable deployment of DNNs, especially in safety-critical systems, since it is difficult to anticipate all potential types of corruption the model might encounter in production.
2.2 Multiplicative weight perturbations simulate input corruptions
Figure 2: Depiction of how a corruption $\mathbf{g}$ affects the output of a DNN. Here $\mathbf{x}_{\mathbf{g}}=\mathbf{g}(\mathbf{x})$. The corruption $\mathbf{g}$ creates a shift $\boldsymbol{\delta}_{\mathbf{g}}\mathbf{x}^{(0)}=\mathbf{x}_{\mathbf{g}}-\mathbf{x}$ in the input $\mathbf{x}$, which propagates into shifts $\boldsymbol{\delta}_{\mathbf{g}}\mathbf{x}^{(h)}$ in the output of each layer. This will eventually cause a shift in the loss $\boldsymbol{\delta}_{\mathbf{g}}\ell$. This figure explains why the model performance tends to degrade under corruption.
To address the problem above, we make two key assumptions about the corruptions in $\mathcal{G}$:
@@ 2.2 Multiplicative weight perturbations simulate input corruptions
A model’s robustness to corruptions in $\mathcal{G}$ can be indirectly enhanced by improving its resilience to a more easily simulated set of input perturbations.
Assumption 1 implies that the corrupted versions of an input $\mathbf{x}$ must be constrained within a bounded neighborhood of $\mathbf{x}$ in the input space. Assumption 2 is corroborated by Rusak et al. (2020), who demonstrated that distorting training images with Gaussian noise improves a DNN’s performance against various types of corruption. We further validate this observation for corruptions beyond Gaussian noise in Section 4.1. However, Section 4.1 also reveals that using corruptions as data augmentation degrades model performance on clean images. Consequently, we need to identify a method that efficiently simulates diverse input corruptions during training, thereby robustifying a DNN against a wide range of corruptions without compromising its performance on clean inputs.
One such method involves injecting random multiplicative weight perturbations (MWPs) into the forward pass of DNNs during training. The intuition behind this approach is illustrated in Fig. 1. Essentially, for a pre-activated neuron $z=\mathbf{w}^{\top}\mathbf{x}$ in a DNN, given a corruption causing a covariate shift $\boldsymbol{\epsilon}$ in the input $\mathbf{x}$, Figs. 1(a) and 1(b) show that one can always find an equivalent MWP $\boldsymbol{\xi}(\boldsymbol{\epsilon},\mathbf{x})$:
$z=\mathbf{w}^{\top}(\mathbf{x}+\boldsymbol{\epsilon})=(\mathbf{w}\circ\boldsymbol{\xi}(\boldsymbol{\epsilon},\mathbf{x}))^{\top}\mathbf{x},\quad\boldsymbol{\xi}(\boldsymbol{\epsilon},\mathbf{x})=1+\boldsymbol{\epsilon}/\mathbf{x}$ (2)
where $\circ$ denotes the Hadamard product. This observation suggests that input corruptions can be simulated during training by injecting random MWPs into the forward pass, as depicted in Fig. 1(c), resulting in a model more robust to corruption. We thus move the problem of simulating corruptions from the input space to the weight space.
Here we provide theoretical arguments supporting the usage of MWPs to robustify DNNs. To this end, we study how corruption affects training loss. We consider a feedforward neural network $\mathbf{f}(\mathbf{x};\boldsymbol{\omega})$ of depth $H$ parameterized by $\boldsymbol{\omega}=\{\mathbf{W}^{(h)}\}_{h=1}^{H}\in\mathcal{W}$, which we define recursively as follows:
$\mathbf{f}^{(0)}(\mathbf{x})\vcentcolon=\mathbf{x},\quad\mathbf{z}^{(h)}(\mathbf{x})\vcentcolon=\mathbf{W}^{(h)}\mathbf{f}^{(h-1)}(\mathbf{x}),\quad\mathbf{f}^{(h)}(\mathbf{x})\vcentcolon=\boldsymbol{\sigma}^{(h)}(\mathbf{z}^{(h)}(\mathbf{x})),\quad\forall h=1,\dots,H$ (3)
where $\mathbf{f}(\mathbf{x};\boldsymbol{\omega})\vcentcolon=\mathbf{f}^{(H)}(\mathbf{x})$ and $\boldsymbol{\sigma}^{(h)}$ is the non-linear activation of layer $h$. For brevity, we use $\mathbf{x}^{(h)}$ and $\mathbf{x}^{(h)}_{\mathbf{g}}$ as shorthand notations for $\mathbf{f}^{(h)}(\mathbf{x})$ and $\mathbf{f}^{(h)}(\mathbf{g}(\mathbf{x}))$ respectively. Given a corruption function $\mathbf{g}$, Fig. 2 shows that $\mathbf{g}$ creates a covariate shift $\boldsymbol{\delta}_{\mathbf{g}}\mathbf{x}^{(0)}\vcentcolon=\mathbf{x}_{\mathbf{g}}^{(0)}-\mathbf{x}^{(0)}$ in the input $\mathbf{x}$ leading to shifts $\boldsymbol{\delta}_{\mathbf{g}}\mathbf{x}^{(h)}\vcentcolon=\mathbf{x}^{(h)}_{\mathbf{g}}-\mathbf{x}^{(h)}$ in the output of each layer. This will eventually cause a shift in the per-sample loss $\boldsymbol{\delta}_{\mathbf{g}}\ell(\boldsymbol{\omega},\mathbf{x},y)\vcentcolon=\ell(\boldsymbol{\omega},\mathbf{x}_{\mathbf{g}},y)-\ell(\boldsymbol{\omega},\mathbf{x},y)$. The following lemma characterizes the connection between $\boldsymbol{\delta}_{\mathbf{g}}\ell(\boldsymbol{\omega},\mathbf{x},y)$ and $\boldsymbol{\delta}_{\mathbf{g}}\mathbf{x}^{(h)}$:
@@
${\color[rgb]{0.63671875,0.140625,0.6796875}\boldsymbol{\delta}_{\mathbf{g}}\ell(\boldsymbol{\omega},\mathbf{x},y)}\leq{\color[rgb]{0.99609375,0.26171875,0.234375}{\left\langle\nabla_{\mathbf{z}^{(h+1)}}\ell(\boldsymbol{\omega},\mathbf{x},y)\otimes\boldsymbol{\delta}_{\mathbf{g}}\mathbf{x}^{(h)},\mathbf{W}^{(h+1)}\right\rangle}_{F}}+{\color[rgb]{0.03515625,0.5859375,0.9375}\frac{C_{\mathbf{g}}^{(h)}(\mathbf{x})}{2}||\mathbf{W}^{(h)}||_{F}^{2}}$ (4)
Here $\otimes$ denotes the outer product of two vectors, $\langle\cdot,\cdot\rangle_{F}$ denotes the Frobenius inner product of two matrices of the same dimension, $||\cdot||_{F}$ is the Frobenius norm, and $\nabla_{\mathbf{z}^{(h)}}\ell(\boldsymbol{\omega},\mathbf{x},y)$ is the Jacobian of the per-sample loss with respect to the pre-activation output $\mathbf{z}^{(h)}(\mathbf{x})$ at layer $h$. To prove Lemma 1, we use Assumption 1 and the following assumption about the loss function:
Assumption 3.
The input gradient of the per-sample loss $\nabla_{\mathbf{x}}\ell(\boldsymbol{\omega},\mathbf{x},y)$ is Lipschitz continuous.
Assumption 3 allows us to define a quadratic bound of the loss function using a second-order Taylor expansion. The proof of Lemma 1 is provided in Appendix A. Using Lemma 1, we prove Theorem 1, which bounds the training loss in the presence of corruptions using the training loss under multiplicative perturbations in the weight space:
@@
For a function $\mathbf{g}:\mathcal{X}\rightarrow\mathcal{X}$ satisfying Assumption 1 and a loss function $\mathcal{L}$ satisfying Assumption 3, there exists $\boldsymbol{\xi}_{\mathbf{g}}\in\mathcal{W}$ and $C_{\mathbf{g}}>0$ such that:
${\color[rgb]{0.63671875,0.140625,0.6796875}\mathcal{L}(\boldsymbol{\omega};\mathbf{g}(\mathcal{S}))}\leq{\color[rgb]{0.99609375,0.26171875,0.234375}\mathcal{L}(\boldsymbol{\omega}\circ\boldsymbol{\xi}_{\mathbf{g}};\mathcal{S})}+{\color[rgb]{0.03515625,0.5859375,0.9375}\frac{C_{\mathbf{g}}}{2}||\boldsymbol{\omega}||_{F}^{2}}$ (5)
We provide the proof of Theorem 1 in Appendix B. This theorem establishes an upper bound for the target loss in Eq. 1:
${\color[rgb]{0.63671875,0.140625,0.6796875}\mathcal{L}(\boldsymbol{\omega};\mathcal{G}(\mathcal{S}))}\leq\mathbb{E}_{\mathbf{g}\sim\mathcal{G}}\left[{\color[rgb]{0.99609375,0.26171875,0.234375}\mathcal{L}(\boldsymbol{\omega}\circ\boldsymbol{\xi}_{\mathbf{g}};\mathcal{S})}+{\color[rgb]{0.03515625,0.5859375,0.9375}\frac{C_{\mathbf{g}}}{2}||\boldsymbol{\omega}||_{F}^{2}}\right]$ (6)
This bound implies that training a DNN using the following loss function:
${\color[rgb]{0.63671875,0.140625,0.6796875}\mathcal{L}_{\boldsymbol{\Xi}}(\boldsymbol{\omega};\mathcal{S})}\vcentcolon=\mathbb{E}_{\boldsymbol{\xi}\sim\boldsymbol{\Xi}}\left[{\color[rgb]{0.99609375,0.26171875,0.234375}\mathcal{L}(\boldsymbol{\omega}\circ\boldsymbol{\xi};\mathcal{S})}\right]+{\color[rgb]{0.03515625,0.5859375,0.9375}\frac{\lambda}{2}||\boldsymbol{\omega}||_{F}^{2}}$ (7)
where the expected loss is taken with respect to a distribution $\boldsymbol{\Xi}$ of random MWPs $\boldsymbol{\xi}$, will minimize the upper bound of the loss $\color[rgb]{0.63671875,0.140625,0.6796875}\mathcal{L}(\boldsymbol{\omega};\hat{\mathcal{G}}(\mathcal{S}))$ of a hypothetical set of corruptions $\hat{\mathcal{G}}$ simulated by $\boldsymbol{\xi}\sim\boldsymbol{\Xi}$. This approach results in a model robust to these simulated corruptions, which, according to Assumption 2, could indirectly improve robustness to corruptions in $\mathcal{G}$.
@@
5:     Draw a mini-batch =$\mathcal{B}=\{(\mathbf{x}_{b},y_{b})\}_{b=1}^{B}\sim\mathcal{S}$.
@@
We note that the second term in Eq. 7 is the $L_{2}$-regularization commonly used in optimizing DNNs. Based on this proxy loss, we propose Algorithm 1 which minimizes the objective function in Eq. 7 when $\boldsymbol{\Xi}$ is an isotropic Gaussian distribution $\mathcal{N}(\boldsymbol{1},\sigma^{2}\mathbf{I})$. We call this algorithm Data Augmentation via Multiplicative Perturbations (DAMP), as it uses random MWPs during training to simulate input corruptions, which can be viewed as data augmentations.
@@
The standard method to calculate the expected loss in Eq. 7, which lacks a closed-form solution, is the Monte Carlo (MC) approximation. However, the training cost of this approach scales linearly with the number of MC samples. To match the training cost of standard SGD, Algorithm 1 divides each data batch into $M$ equal-sized sub-batches (6) and calculates the loss on each sub-batch with different multiplicative noises from the noise distribution $\boldsymbol{\Xi}$ (79). The final gradient is obtained by averaging the sub-batch gradients (11). Algorithm 1 is thus suitable for data parallelism in multi-GPU training, where the data batch is evenly distributed across $M>1$ GPUs. Compared to SGD, Algorithm 1 requires only two additional operations: generating Gaussian samples and point-wise multiplication, both of which have negligible computational costs. In our experiments, we found that both SGD and DAMP had similar training times.
@@
In this section, we demonstrate that optimizing DNNs with adversarial MWPs follows a similar update rule to Adaptive Sharpness-Aware Minimization (ASAM) (Kwon et al., 2021). We first provide a brief description of ASAM and its predecessor Sharpness-Aware Minimization (SAM) (Foret et al., 2021):